OpenStudy (anonymous):

H8P3: MEMORY Solutions anyone?

OpenStudy (anonymous):

185.34 5 3.8 3RD---(-ln( (3.5-5) / (0.5-5) ) * (RPU+RON) * 3.5*10^-15) * 10^12 4TH----(-ln( (0.9-(5*1800/(RON+RPU))) / (3.5-(5*RON/(RON+RPU))) ) * ((RON|| RPU) + RON) * 3.5*10^-15) * 10^12 If u got correct give me best response :D

OpenStudy (anonymous):

We assume an "ancient" 1μm technology satisfying the static discipline: VS=5.0V, VOH=3.5V, VIH=3.0V, VIL=0.9V, VOL=0.5V The gate-source capacitance CGS=3.5fF and VT=1.0V. RON=2150.0Ω, ROFF=115.0MΩ and RPU=12.0kΩ. Can you please solve the last 2 questions for these values . Cuz I dint get right answers :(

OpenStudy (anonymous):

H8P1 Impulse L=5mH, R1=10Ω, C=33nF and R2=10kΩ. (a) What is the value of vR1(t) at t=0− in Volts (V)? (b) What is the value of vR1(t) at t=0+ in volts (V)? Hint: Recall that that ∫0+0−δ(t)dt=1 (c) What is the value for vR1(t) at t=1ms in volts(V)? (Hint: The L-R1 and C-R2 branches of the circuit are decoupled.) ? ?

OpenStudy (anonymous):

3rd Answer- 46.5262304251